OpenStudy (anonymous):

If f(x) = -x² - 4x + 3, find f(4)

OpenStudy (anonymous):

Plug in 4 for x.

OpenStudy (zzr0ck3r):

you have -(x)^2-4(x)+3 put 4 where you see an x

OpenStudy (zzr0ck3r):

note:-(x)^2 does not equal (-x)^2 and you have the fist one, so that term -(x)^2 when x =4 is -(4)^2 = -16

OpenStudy (anonymous):

4x^3-17x^2-42x/x^2y-13xy+42y

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

thats what im sayin lol

jhonyy9 (jhonyy9):

what you wann to do with ?

OpenStudy (anonymous):

simplify

jhonyy9 (jhonyy9):

have you idea for ?

OpenStudy (jhannybean):

\[\large f(x) = -x^{2} - 4x + 3\]Then \[\large f(4) = -(4)^{2} - 4(4) + 3\]\[\large f(4) =?\]

jhonyy9 (jhonyy9):

4x^3-17x^2-42x ---------------- = ? x^2y-13xy+42y

OpenStudy (anonymous):

it doesnt give an =